# ChatWise: AI-Powered Engaging Conversations for Enhancing Senior Cognitive Wellbeing

### v1

@@ 4.1 Conversation Generation
For controlled and responsive development of ChatWise, we employed 9 digital twins provided by  Hong et al. (2024) as simulated users with different personas, which are fine-tuned LLMs to approximate senior individuals with MCI symptoms using real, de-identified dialogue data from a clinical trial  I-CONECT (2024). We collected 20 trajectories of conversations between ChatWise and each digital twin, where a conversation contains 10 turns, each turn with a two-way utterance exchange between ChatWise and digital twin (user).
4.2 Engagement Evaluation
@@
User verbosity: We primarily measure user engagement through verbosity, i.e. the user-to-chatbot token ratio, defined as $\frac{\text{tokens}_{user}}{\text{tokens}_{chatbot}}$, as a clinical study (Yu et al., 2021) suggested reducing the moderator talkativeness while encouraging participant (user) expression.
@@
Cognitive Metrics: See et al. (2019) defines 9 aspects for evaluating conversation quality, from which we select 5 focused on engagement assessment: Engagingness, Interestingness, Listening, Fluency, and Making Sense. While Interestingness assesses overall dialogue quality, the others primarily evaluate user interactions.
We compute the win rate (Rafailov et al., 2023) by comparing dialogues generated with and without ChatWise, using the same pretrained LLM (GPT-4o) as the utterance generator and identical system prompts, except for input from $\pi_{s}$ to $\pi_{u}$. Dialogue pairs are randomly matched within the same digital twin for evaluation.
@@
Strategy Provider Verbosity $\uparrow$ Engagingness WR $\uparrow$ Interestingness WR $\uparrow$ Listening WR $\uparrow$ Fluency WR $\uparrow$ Making sense WR $\uparrow$
w/o ChatWise 0.7398 0.3389 0.3833 0.3593 0.4111 0.4056
GPT-4o 0.8635 0.6833 0.6222 0.6222 0.5778 0.5833
o3-mini 0.8643 0.6778 0.6222 0.6778 0.6222 0.6167
Llama 3.1-405B 0.8083 0.6222 0.6056 0.6222 0.5667 0.5833
Table 1: Performance on 10-turn dialogue across different strategy providers. W/o ChatWise is the baseline without applying any strategy, whose win rates (WRs) are averaged across strategy providers. (Appendix A.3)
4.3 ChatWise Performance and Analysis
We evaluated ChatWise using different LLM backbones as strategy providers, including GPT-4o, o3-mini, and Llama3.1-405B. As shown in Table 1, ChatWise consistently outperformed non-strategic dialogues across all evaluation metrics.
ChatWise robustly enhances user engagement: Regardless of the capability of pretrained LLMs, our design consistently improves user engagement and dialogue quality with a large margin across all tested LLMs as the strategy provider.
@@ 4.3 ChatWise Performance and Analysis
Reasoning ability can transfer to improve engagement: The best-performing model in our setting, o3-mini, was optimized for STEM reasoning tasks, which indicates that strong reasoning ability can also benefit inferring dialogue strategies.
4.4 Accumulated Multi-Turn Analysis
We analyzed ChatWise’s performance over increasing conversation turns, as shown in Figure 3, which presents accumulated metrics and win rates. Results revealed a clear upward trend in all key engagement metrics—especially verbosity, engagingness, interestingness, and listening—demonstrating that ChatWise’s effects strengthen as dialogues progress, while baseline gradually degrades, which highlights ChatWise’s capability in enhancing long-term conversational engagement.
(a) Verbosity.
(b) Engagingness Win Rate.
(c) Interestingness Win Rate.
(d) Listening Win Rate.
Figure 3: ChatWise performance with o3-mini as strategy provider. W/o ChatWise is the baseline and GPT-4o servers as utterance generator in all settings. Evaluation starts by skipping two warming-up turns (Sec 3).
@@ 4.5 Ablation Study
We evaluated ChatWise without user emotion in the strategy provider’s output (Table 2). Its consistent performance confirms that strategies are the primary driver of engaging conversations.
Strategy Provider GPT-4o GPT-4o w/o emotion w/o ChatWise
@@
Table 2: User verbosity under different strategy providers. GPT-4o w/o emotion keeps the strategy provider with user emotion information removed.
@@ 4.6 ChatWise’s Robustness to User Persona
Figure 4 shows the significant gain of ChatWise interacting with all simulated users, indicating its robustness to support varying senior characteristics.
Figure 4: Users’ dialogue verbosity (log-normalized, Appendix A.5) across 9 digital twins using different LLMs as strategy providers, compared with baseline (w/o ChatWise).
4.7 Dialogue Analysis
We identified 4 digital twins for which ChatWise was most effective and collected 40 additional dialogues from each to conduct a deeper analysis, with the main findings summarized below.
Transient user emotions: We defined emotion transition triplets as $(\text{e}_{i},\text{s},\text{e}_{i+1})$, where $\text{e}_{i}$ ($\text{e}_{i+1}$) represents the user’s emotion at turn $i$ ($i+1$), and $s$ is the strategy provided by ChatWise at turn $i$. We computed the average occurrence of emotion triplets and showed the top 15 most frequent in Figure 5. Most triplets reflect unchanged user emotions, indicating the challenge of detecting or influencing emotional shifts within a short turn.
Long-term user emotions: We analyzed user emotion changes from the beginning to the end of dialogues and calculated averaged occurrence across digital twins (Figure 6). Over 48% users experienced emotional shifts post-dialogue, with significantly more positive changes after engaging with ChatWise. This implies that while transient emotions are difficult to track, strategic conversational support can improve user emotions over time.
Predominant strategies: We identified the top 10 most frequent strategies across digital twins and found that predominant interaction strategies remained consistent across user types (See Appendix A.8). Particularly, Open Question, Statement-non-opinion, and Acknowledgment strategies dominate ChatWise driven conversations, suggesting their potential efficacy in fostering engaging dialogues.
@@ 4.7 Dialogue Analysis
Figure 6: Occurrence of user emotional changes. For instance, “neutral-joy” indicates user emotion transitioning from neutral (conversation begins) to joy (ends).
@@ 5 Conclusion and Future Work
We introduce ChatWise, a tuning-free yet principled framework for developing AI chatbots that foster long-term user engagement. Our dual-level generation design integrates clinical insights into LLM reasoning and is compatible with various LLMs. Extensive evaluations based on de-identified clinical data showed that our method robustly enhances the engagement and cognitive status of seniors. Moving forward, we aim to 1) enhance multi-turn interaction via RL and 2) validate its effectiveness through real user studies.
@@
All digital twins were provided as fine-tuned GPT-3.5 APIs, and its high economic cost, at the time when this project was going, constrained the volume of dialogues we generated. To mitigate this, future work will explore training digital twins using LLaMA 3.1 to reduce sampling costs while maintaining conversational quality. Additionally, our study relies on simulated dialogues rather than real user interactions. A real user study is needed to validate the system’s effectiveness in dynamic settings, which we plan to incorporate into our future research.
@@ A.1 Prompt design
The prompt used in the experiments and structured output design are available at https://anonymous.4open.science/r/ChatWise-D5E0, including:
@@
A.2 Data Generation Configuration
We used GPT-4o as the utterance generator in ChatWise and tested different LLM backbones as strategy providers. Considering the Llama3.1-405B does not support structured output, we applied GPT-4o as the strategy extractor to structure the strategy output of Llama3.1-405B. The following are the configurations of each model:
@@ A.2 Data Generation Configuration
A.3 Win rate for w/o ChatWise
@@ A.4 Dialogue Acts
Table 3: Dialogue Act to its corresponding tag.
@@
A.5 Log-normalization
$y=\ln{\left(4x+1\right)}$
A.6 Additional results for Accumulated Multi-Turn Analysis
ChatWise’s accumulated performance on Fluency, and Making Sense is shown in Figure 7.
(a) Fluency Across Turns
(b) Making sense Across Turns
Figure 7: ChatWise performance with o3-mini as strategy provider across turns. W/o ChatWise is the baseline that ChatWise is not applied. GPT-4o servers as moderator in all settings.
A.7 Primary Strategies
Figure 8: Strategy occurrence across digital twins.

### v2

@@ 4.2.1 Conversation Generation
While evaluation on real clinic data shows strong alignment between ChatWise and professional caregivers, the conversations are fixed, thus lacking indications whether and how ChatWise may influence user cognitive or emotional status during conversations. To enable controlled and responsive development of ChatWise prior to future human user studies, we used 9 digital twins provided by  Hong et al. (2024) as simulated users with different personas, which are fine-tuned LLMs calibrated to mimic the linguistic behaviors of older adults, some of whom show MCI symptoms. All digital twins are trained using real, de-identified dialogue data from the I-CONECT clinical trial. We collected 20 trajectories of conversations between ChatWise (moderator) and each digital twin as a user, with each conversation containing 20 turns. The same settings were applied to the baseline chatbot for fair comparisons.
4.2.2 Metrics for Interactive Engagement
@@
User Verbosity: Given a conversation sequence $\tau$ that consists of user utterance $x$ and moderator utterance $u$, we primarily measure user engagement through the user’s talkativeness compared to the moderator, i.e. the user’s verbosity ($v$), defined as $v=\frac{\sum_{x\in\tau}|x|}{\sum_{u\in\tau}|u|},$ where $|x|$ ($|u|$) denotes the number of tokens in an utterance $x$ ($u$). This metric follows a clinical study (Yu et al., 2021) that suggests reducing the moderator talkativeness while encouraging participant (user) expression.
Cognitive Win Rate:
@@
See et al. (2019) defined different aspects for evaluating conversation quality, from which we select 3 that are focused on assessing user cognitive status: Listening, Fluency, and Making Sense. All three metrics evaluate the user behavior rather than the overall dialogue quality considering both user and moderator behaviors.
Instead of providing a numerical score for each cognitive metric that might be unstable, we adopt the Win Rate (WR) definition from prior work  (Rafailov et al., 2023) to compare pairs of dialogues generated with and without ChatWise. Dialogue pairs are randomly matched so long as they are collected from the same digital twin (simulated user) for evaluation. WR is defined as the proportion of pairs in which the dialogue generated with ChatWise is preferred over the baseline using an LLM-as-judge, based on a given cognitive metric.
To mitigate the LLM’s position bias issues observed in prior work Shi et al. (2024); Yu et al. (2024), we excluded sample pairs with inconsistent preference labels when the order of the two samples was reversed and calculated WRs using only the remaining consistent pairs.
@@
Strategy Provider Verbosity $\uparrow$ Listening $\uparrow$ Fluency $\uparrow$ Making Sense $\uparrow$
baseline 0.7398 0.5249 0.4986 0.5024
GPT-4o 0.8635 0.4962 0.4748 0.4786
o3-mini 0.8643 0.4884 0.5368 0.5180
Llama 3.1-405B 0.8083 0.4407 0.4926 0.4963
Table 1: Multifaceted evaluations on cognitive engagement with data sampled by interacting with digital twins over 20 turns, using different LLMs serving as strategy providers, where Listening, Fluency, and Making Sense metrics report win rates, and baseline denotes a conversation generator without receiving strategy guidance. All evaluations focus on user reactions.
4.2.3 Models and Baseline
We evaluated ChatWise with 3 different LLM backbones as strategy providers: GPT-4o, o3-mini, and Llama3.1-405B. For each setting, a baseline adopts the same utterance generator as ChatWise while without receiving strategy guidance. The WR of baseline is reported as the complement of ChatWise’s average WR: $\text{WR}_{\text{baseline}}=1-\text{average}(\text{WR}_{GPT-4o}+\text{WR}_{o3-mini}+\text{WR}_{Llama3.1-405B}),$ where $\text{WR}_{X}$ indicates the WR of ChatWise, given model $X$ as the strategy provider.
4.2.4 Performance Overview
Effectiveness of Strategy-Guided Generation:
@@ 4.2.4 Performance Overview
As shown in Table 1, our design mostly demonstrated better user engagement compared to baselines across tested LLMs as the strategy provider. Particularly, the best-performing model in our setting, o3-mini, was optimized for STEM reasoning tasks, which indicates that strong reasoning ability can also benefit inferring dialogue strategies.
(a) User Verbosity.
(b) WR: Listening.
(c) WR: User Fluency.
(d) WR: Making sense.
Figure 8: Multifaceted Engagement Evaluation across conversation turns, excluding the first two warm-up turns. ChatWise’s performance gain over the baseline exists over multi-turn dialogues.
Performance with Dialogue Progression:
We analyzed ChatWise’s performance over increasing dialogue length by truncating conversations at turn $t$ and computing metrics accordingly, as shown in Figure 8. Results revealed a clear upward trend in verbosity, which indicates increased talkativeness of the user rather than the ChatWise, as opposed to the baseline. Meanwhile, the WR gains of ChatWise either remain stable or show a slight decline as conversations progress, which we infer are subject to the length bias of LLMs Dubois et al. (2024) as the overall number of tokens of ChatWise-generated dialogues tends to decrease in later turns.
@@ 4.3 Ablation Study
We evaluated ChatWise with and without user emotion in the strategy provider’s output. As shown in Table 2, where both strategy providers used GPT-4o as the backbone, removing emotion information led to only a slight drop of performance, while removing ChatWise entirely resulted in a much lower verbosity score, which highlights that strategy guidance itself is the key driver of user engagement.
Method ChatWise ChatWise w/o emotion w/o ChatWise
@@
Table 2: Ablative study on the performance of user verbosity when removing different contextualized information. w/o emotion denotes the strategy provider without user emotion information as input.
@@ 4.4 ChatWise’s Robustness to User Persona
As shown in Figure 9, the gain of ChatWise interacting with 9 simulated users is consistently significant, in which the metrics were log-normalized. This indicates the robustness of ChatWise  given varying senior characteristics.
Figure 9: User verbosity log-normalized across 9 digital twins. Given varying strategy provider backbones, our design consistently outperforms the baseline (w/o ChatWise) with non-strategic conversations.
4.5 Analysis of User Emotion Transitions
To analyze if ChatWise can support positive emotional shifts during conversations, we selected four digital twins and collected 40 additional dialogues from each for deeper analysis, with the main findings summarized below.
Transient user emotions:
We define an emotion transition triplet as $(\text{e}_{t},u_{t},\text{e}_{t+1})$, where $\text{e}_{t}$ ($\text{e}_{t+1}$) represents the user’s emotion at turn $t$ ($t+1$), and $u_{t}$ is the moderator utterance in between. We computed the average occurrence of emotion triplets and showed the top 15 most frequent in Figure 10. Most triplets reflect unchanged user emotions, which indicates difficulty in either influencing or detecting emotional shifts within a short turn.
@@ 4.5 Analysis of User Emotion Transitions
User Emotions over Multi-Turn Conversations:
We then analyzed user emotion changes from the beginning to the end of dialogues and calculated averaged occurrence across digital twins (Figure 11(b)). Over 48% users experienced emotional shifts post-dialogue, with significantly more positive changes after engaging with ChatWise. This implies that while transient emotions are difficult to track, strategic conversational support may improve user emotions over multi-turn conversations.
(a) Real user emotion transitions summarized from 150 dialogues in the I-CONECT study.
(b) Digital twin emotion transitions summarized over 80 dialgoues.
Figure 11: User emotion shifts in conversations between digital-twin and ChatWise (Figure 11(b)) align well with real-user emotion shifts in the clinical study (Figure 11(a)). More neutural-to-joy transitions were observed given ChatWise compared to the baseline.
Predominant Strategies:
We identified the top 10 most frequent strategies across digital twins (Figure 13) and found that predominant interaction strategies remained consistent across user types (See Appendix A.8). Particularly, Open Question, Statement-non-opinion, and Acknowledgment strategies dominate ChatWise driven conversations.
@@ 5 Conclusion and Future Work
We present a dual-level framework for AI chatbots that supports multi-turn conversations for older adults by integrating clinical insights into LLM reasoning. Extensive evaluations based on both real clinical data and generated conversations showed that our method aligns well with the behavior of professional human caregivers while robustly enhancing the cognitive status of older adults through simulation studies. Our future work includes training-based methods for further optimizing multi-turn interaction experience and investigating the acceptability and feasibility of our design through real user studies.
6 Ethical Considerations
This study involved simulated dialogues using digital twins trained on de-identified conversation data from a clinical trial involving older adults. Offline analysis of this dataset was conducted to evaluate the similarity between chatbot and caregiver strategies. This project did not use any personally identifiable information, and all data processing followed institutional privacy and research ethics guidelines. Potential risks are controllable, which include privacy concerns of emotion detection and misinterpretation of conversation due to AI-based evaluation. To mitigate this, the emotional detection module in the system is made optional and can be omitted as configured. Win rate calculations are based on a sanity check where we omitted pair samples that receive inconsistent preference labels by LLM-as-the-judge after swapping their positions. We have manually cross-validated a sufficient subset of AI-generated analyses to reduce bias. A real-user study is planned for future validation and will undergo full Institutional Review Board (IRB) approval.
@@ 6 Ethical Considerations
All digital twins were provided as fine-tuned GPT-3.5 APIs, and their high economic cost, at the time when this project was going on, constrained the volume of dialogues we generated. To mitigate this, future work will explore accessing digital twins using LLaMA 3.1 to reduce sampling costs while maintaining conversational quality. LLM evaluations have been discussed to show position bias. Regarding win rate calculation, after we took mitigation methods and filtered out pairs with inconsistent preference labels, the remaining samples contributed to 81% of the data before filtering. Future mitigation methods will include learning a reward function for more fine-grained and robust evaluation. Furthermore, our study relies either on offline static data or simulated dialogues rather than dynamic user interactions. A real user study is needed to validate the system’s acceptability and feasibility in real-world settings, which we plan to incorporate into our future research.
@@ A.1 Prompt design
The prompt used in the experiments and structured output design are available at https://anonymous.4open.science/r/ChatWise-8F53, including:
@@
A.2 Offline Data Preparation
We extracted dialogue content from the I-CONECT video conversations and processed it into OpenAI-compatible dialogue history format. Dialogues with fewer than 40 turns were excluded. We then subsampled two subsets for evaluation purposes. To assess the overall strategy alignment of ChatWise, we randomly sampled 150 dialogues. To evaluate the robustness of ChatWise across different participants and time periods, we randomly selected 7 participants who completed the full study and sampled one dialogue from each of their sessions during weeks 1, 9, 17, 25, 33, and 41. All user identifiers and personally identifiable information have been anonymized.
The OpenAI-compatible dialogue history format example:[{"role": "system", "content": system_prompt}, {"role": "user", "content": user_content1}, {"role": "assistant", "content":assistant_content1}, {"role": "user", "content": user_content2}, {"role": "assistant", "content": assistant_content2}, …].
A.3 Data Generation Configuration
By default, we used o3-mini as strategy provider. GPT-4o serves as an utterance generator in all settings. W/o ChatWise denotes the baseline, which is using GPT-4o as the utterance generator only, without a strategy provider. We tested different LLM backbones as strategy providers in ChatWise. Considering the Llama3.1-405B does not support structured output, we applied GPT-4o as the strategy extractor to structure the strategy output of Llama3.1-405B. We employed GPT-4o as the judge to select the preferred one when computing the Win Rate. The prompt for selecting the preferred response is listed in Figure 12. The following are the configurations of each model:
@@ A.3 Data Generation Configuration
GPT-4o as judge:
n=1
max_tokens=1024
top_p=1
temperature=1
Figure 12: Prompt for GPT-4o as a Judge.
A.4 Win rate for w/o ChatWise
@@ A.5 Dialogue Acts
A.6 Log-normalization
$y=\ln{\left(4x+1\right)}$
A.7 Primary Strategies
Figure 13: Strategy occurrence across digital twins.
@@ A.7 Primary Strategies
Table 3: Strategy to its corresponding tag. The strategies are drawn from the DAs in DAMSL  (Allen and Core, 1997) that are used by telehealth clinical trials  (Yuan et al., 2023), integrated strategies from prior emotional support dataset  (Liu et al., 2021) .